# Predictive Control Barrier Functions: Bridging model predictive control and control barrier functions

### v1

@@ Abstract
In this paper, we establish a connection between model predictive control (MPC) techniques and Control Barrier Functions (CBFs). Recognizing the similarity between CBFs and Control Lyapunov Functions (CLFs), we propose a safe MPC formulation that ensures invariance and safety without relying on explicit stability conditions. The value function of our proposed safe MPC is a CBF, which we refer to as the Predictive Control Barrier Function (PCBF), similar to traditional MPC formulations which encode stability by having value functions as CLFs. Our formulation is simpler than previous PCBF approaches and is based on weaker assumptions while proving a similar theorem that guarantees safety recovery. Notably, our safe MPC formulation does not require the value function to be strictly decreasing to ensure convergence to a safe invariant set. Numerical examples demonstrate the effectiveness of our approach in guaranteeing safety and constructing non-conservative CBFs.
@@ I Introduction
The development of new control methods that can provide strong safety guarantees is critical due to the increasing prevalence of autonomy in safety-critical applications. The challenge of safety is that systems under unsafe control actions may not immediately violate state constraints, but can nevertheless inevitably reach unsafe regions. Safety can therefore be guaranteed by ensuring positive invariance of the safe set [1]. A set is positive invariant if the inclusion of the state at some times implies the inclusion in the future [2].
@@
In this work we propose a safe MPC formulation to guarantee positive invariance and safety. It constructs a non-conservative safe invariant set and ensures convergence of trajectories that start from initially unsafe regions to the maximal safe invariant set. In contrast to traditional MPC convergence proofs that rely on Lyapunov’s theorem [14], requiring strictly decreasing value functions, we apply the invariant set theorem [16, Section 3.4.3] to guarantee convergence with weaker conditions on a non-increasing value function.
@@
  • We apply our MPC formulation to linear and nonlinear examples, demonstrating its ability to expand a predefined conservative safe invariant set and guarantee local convergence of trajectories to a safe invariant set.
@@ III Safe MPC
We next present our safe MPC formulation with soft state constraints. We show that all trajectories, starting from the feasible set of the safe MPC formulation, remain in or converge to the maximal safe invariant set of the system. The safe MPC problem is formulated as
@@
where $\xi_{i|k}\in\mathbb{R}_{\geq 0}$ is a slack variable for the state constraints $i$ steps ahead when evaluated at the current time $k$, $\mathbb{X}(\xi_{i|k})$ is the relaxed state constraint set, where $\mathbb{X}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}:c(x)\leq\xi_{i|k}\}$. The terminal set $\mathbb{X}_{f}$ is compact and $\mathbb{X}_{f}\subset\mathbb{X}$. We construct $\mathbb{X}_{f}$ such that there exists a local controller $\kappa_{f}(\cdot):\mathbb{X}_{f}\rightarrow\mathbb{U}$, $\forall x\in\mathbb{X}_{f}$, which ensures the positively invariance of $\mathbb{X}_{f}$, i.e., $f(x,\kappa_{f}(x))\in\mathbb{X}_{f},\forall x\in\mathbb{X}_{f}$. We use $\mathbb{X}^{\xi}_{MPC}$ to denote the feasible set of (8). The feasible set of the same MPC problem without constraint relaxation, i.e., let $\xi_{i|k}=0$ for all $i$, is denoted by $\mathbb{X}_{MPC}^{0}$. We assume $\mathbb{X}_{MPC}^{0}$ to have a non-empty interior. The relationship between the different sets is shown in Fig 1.
@@ III-A Convergence of the MPC formulation
Proof. Denote the optimal input, state and constraint violation sequence [^1] In case of multiple minimizers, a solution $\textbf{u}^{*}_{k}$ can be singled out by the application of consistent tie-break rules, e.g., minimizing the $l2$-norm of $\textbf{u}^{*}_{k}$. for problem (8) at current time $k$ by
@@
where $u^{*}_{i|k}\in\mathbb{U},\;\xi^{*}_{i|k}\geq 0,\;\forall i\in\{0,1,...,N-1\}$ and $x^{*}_{N|k}\in\mathbb{X}_{f}$. The corresponding value function is
@@ IV Numerical Examples
This section presents two examples illustrating the efficacy of our results. The first demonstrates the convergence property of the proposed MPC formulation using an unstable linear system, showing that the system can recover from initially unsafe states to safe states. The second example, which is nonlinear, compares the safe invariant set of PCBFs with competing methods. In both examples, we use the MPT3 toolbox [23] to compute maximal invariant sets. In the linear example, the MPC problem is solved using CVXPY [24] with the Clarabel solver [25]. In the nonlinear example, we use the Casadi framework [26] with the IPOPT solver [27] to solve the nonlinear MPC problem.
@@ IV-B Nonlinear Example
Figure 4: Compare the safe invariant sets derived from various methods. The predictive safe set of the PCBF, the maximal invariant set, and the safe invariant set of the hand-crafted CBF are represented by the black line, grey region, and blue line, respectively. It also shows the contours of the PCBF $V^{*}$.

### v2

@@ Abstract
In this paper, we establish a connection between model predictive control (MPC) techniques and Control Barrier Functions (CBFs). Recognizing the similarity between CBFs and Control Lyapunov Functions (CLFs), we propose a MPC formulation that ensures invariance and safety without relying on explicit stability conditions. The value function of our proposed MPC is a CBF, which we refer to as the Predictive Control Barrier Function (PCBF), similar to traditional MPC formulations which encode stability by having value functions as CLFs. Our formulation is simpler than previous PCBF approaches and is based on weaker assumptions while proving a similar theorem that guarantees safety recovery. Notably, our MPC formulation does not require the value function to be strictly decreasing to ensure convergence to a safe invariant set. Numerical examples demonstrate the effectiveness of our approach in guaranteeing safety and constructing non-conservative CBFs.
@@ I Introduction
The development of new control methods that can provide safety guarantees is critical due to the increasing prevalence of autonomy in safety-critical applications. The challenge of safety is that systems under unsafe control actions may not immediately violate state constraints, but can nevertheless inevitably reach unsafe regions. Safety can therefore be guaranteed by ensuring positive invariance of the safe set [1]. A set is positive invariant if the inclusion of the state at some times implies the inclusion in the future [2].
@@
In this work we propose a safe MPC formulation to guarantee positive invariance and safety. It constructs a non-conservative safe invariant set and ensures convergence of trajectories that start from initially unsafe regions to the maximal safe invariant set. In contrast to traditional MPC convergence proofs that rely on Lyapunov’s theorem [14], requiring strictly decreasing value functions, we apply the local invariant set theorem [16, Section 3.4.3] to guarantee convergence with weaker conditions on a non-increasing value function.
@@
  • We apply our MPC formulation to linear and nonlinear examples, demonstrating its ability to expand a predef-ined conservative safe invariant set and guarantee local convergence of trajectories to a safe invariant set.
@@ III Safe MPC
We next present our safe MPC formulation with soft state constraints [^1] While our formulation relaxes all state constraints for simplicity, the theoretical guarantees extend to cases where only partial state constraint relaxation is applied, with the remaining state constraints enforced as hard constraints over the prediction horizon.. We show that all trajectories, starting from the feasible set of the safe MPC formulation, remain in or converge to the maximal safe invariant set of the system. The safe MPC problem is formulated as
@@
where $\xi_{i|k}\in\mathbb{R}_{\geq 0}$ is a slack variable for the state constraints $i$ steps ahead when evaluated at the current time $k$, $\mathbb{X}(\xi_{i|k})$ is the relaxed state constraint set, where $\mathbb{X}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}:c(x)\leq\xi_{i|k}\}$ [^2] To relax state constraints elementwise, we may alternatively employ a slack variable vector $\xi_{i|k}\in\mathbb{R}^{n_{x}}_{\geq 0}$, defining the constrained set as $\mathbb{X}(\xi_{i|k}):=\{x\in\mathbb{R}^{n}:c(x)\leq\xi_{i|k}\}$. This would lead to an objective function of the form $\sum_{i=0}^{N-1}||\xi_{i|k}||$. For notational simplicity, we maintain our current formulation, which effectively corresponds to using a slack vector with the $l_{\infty}$-norm penalty. All theoretical results remain valid when using slack vector formulations.. The terminal set $\mathbb{X}_{f}$ is compact and $\mathbb{X}_{f}\subset\mathbb{X}$. We construct $\mathbb{X}_{f}$ such that there exists a local controller $\kappa_{f}(\cdot):\mathbb{X}_{f}\rightarrow\mathbb{U}$, $\forall x\in\mathbb{X}_{f}$, which ensures the positively invariance of $\mathbb{X}_{f}$, i.e., $f(x,\kappa_{f}(x))\in\mathbb{X}_{f},\forall x\in\mathbb{X}_{f}$. We use $\mathbb{X}^{\xi}_{MPC}$ to denote the feasible set of (8). The feasible set of the same MPC problem without constraint relaxation, i.e., let $\xi_{i|k}=0$ for all $i$, is denoted by $\mathbb{X}_{MPC}^{0}$. We assume $\mathbb{X}_{MPC}^{0}$ to have a non-empty interior. The relationship between the different sets is shown in Fig 1.
@@ III-A Convergence of the MPC formulation
Proof. Denote the optimal input, state and constraint violation sequence [^3] In case of multiple minimizers, a solution $\textbf{u}^{*}_{k}$ can be singled out by the application of consistent tie-break rules, e.g., minimizing the $l_{2}$-norm of $\textbf{u}^{*}_{k}$. for problem (8) at current time $k$ by
@@
where $u^{*}_{i|k}\in\mathbb{U},\;\xi^{*}_{i|k}\geq 0,x^{*}_{i|k}\in\mathbb{X}(\xi^{*}_{i|k}),\;\forall i\in\{0,1,...,N-1\}$ and $x^{*}_{N|k}\in\mathbb{X}_{f}$. The corresponding value function is
@@ IV Numerical Examples
This section presents two examples [^5] While we use 2D examples for illustrative purpose, our safe MPC formulation naturally extends to higher-dimensional problems, like standard MPC approaches. illustrating the efficacy of our results. The first demonstrates the convergence property of the proposed MPC formulation using an unstable linear system, showing that the system can recover from initially unsafe states to safe states. The second example, which is nonlinear, compares the safe invariant set of PCBFs with competing methods. In both examples, we use the MPT3 toolbox [23] to compute maximal invariant sets. In the linear example, the MPC problem is solved using CVXPY [24] with the Clarabel solver [25]. In the nonlinear example, we use the Casadi framework [26] with the IPOPT solver [27] to solve the nonlinear MPC problem.
@@ IV-B Nonlinear Example
Figure 4: Compare the safe invariant sets derived from various methods. The predictive safe set of the PCBF, the maximal invariant set, and the safe invariant set of the hand-crafted CBF are represented by the black line, grey region, and blue line, respectively. It also shows the contours of PCBF $V^{*}$.